# Meta-RL Induces Exploration in Language Agents

### v1

@@ Appendix A Task Description and Details
MineSweeper. We include the classic video game MineSweeper as a partially-observable environment. The game is a 2D square board, with several mines randomly scattered in the board cells. The goal of the game is to open all the safe cells without revealing the hidden mines. In each step, the agent opens a cell, and the first step is always safe. If a mine is revealed, the task ends in failure immediately. The state of the opened (safe) cells can either be empty or a number from 1 to 8, and the number specifies how many mines are adjacent to the specific cell. The agent needs to use the numbers marked on the opened cells to reason about the position of mines. Success is achieved when all safe cells are revealed. Our implementation is based on a simplified version of Li.2023.Minesweeper. The difficulty of this task is controlled by the board size and the number of mines. We train on a board size $6\times 6$ with $3$ hidden mines.
Webshop (yao2022webshop). We include Webshop as a partially-observable text-based environment that simulates online shopping. The agent is given a natural language instruction specifying a product to purchase with certain attributes. The environment presents a simplified e-commerce interface where the agent can search for products, navigate through search results, and examine product pages with details like price, color, size, and customer reviews. The agent must interpret the instruction, search effectively, filter through multiple product options, and select the item that best matches the specified criteria. Success is measured by whether the final purchased item satisfies all the requirements in the original instruction.
ALFWorld (shridhar2020alfworld). We include ALFWorld as a partially-observable text-based environment that simulates household tasks in interactive fiction format. The agent receives natural language instructions for common household activities. The environment provides text descriptions of rooms, objects, and possible actions, while the agent must navigate through a house, interact with objects, and complete multi-step tasks. Objects may need to be found, picked up, cleaned, heated, or combined with other objects to achieve the goal. The agent’s view is limited to the current room and nearby objects, requiring exploration and memory of previously visited locations. Success requires understanding the instruction, planning a sequence of actions, and executing them correctly while managing partial observability. We train on the training examples of the activities ‘Pick’, ‘Look’, ‘Clean’, ‘Heat’. We evaluate in-distribution on the test examples from the same activities, and we evaluate out-of-distribution on the test examples from ‘Cool’, ’Pick2’ activities.
@@ Appendix B Example Prompts
Similar to (feng2025group), we use <action> </action> block to indicate the final decision of the action, and we use <remark> </remark> to indicate the content of the self-reflection.
@@ Appendix C Training details
LaMer is compatible with standard policy gradient algorithms. Without specification, we use GiGPO as the default optimization algorithm. During training, the self-reflection step is also explicitly trained using the reward in the subsequent episodes. During training, we match the total number of experiences sampled for each example between RL and Meta-RL to ensure a fair comparison. Specifically, for each sample $N=3$ episodes and set group size to 8 for Meta-RL, and use a group size of 24 for standard RL training. Besides that, other hyper-parameters and configuration are kept the same between RL and Meta-RL training. We use Qwen3-4B as the base model and train it with Adam optimizer and a learning rate of 1e61e-6. For Sokoban and MineSweeper, we train the agents with a batch size of $16$ for $300$ epochs. In comparison, we use batch size of $8$ and $150$ epochs for Webshop and ALFWorld. The environment reward is set to be $10$ for successful trajectories and 0 for unsuccessful ones. We use temperature of $1.0$ during rollout and $0.7$ during evaluation. The maximum number of output tokens is set to $1024$. Our code is based on the training framework of verl (sheng2025hybridflow) and verl-agent (feng2025group).
@@ D.1 Experiments on different architecture
LaMer is designed as a general framework built on meta-RL principles and is model-agnostic. To validate this, we additionally ran experiments on Llama3.1-8B-Instruct (grattafiori2024llama3herdmodels), showing that our method works well on models in a different model architecture and model size.
@@
In Table 4, we compare LaMer with the strongest baseline GiGPO (feng2025group), on Sokoban and MineSweeper. We show The results indicate that on Llama-3.1-8B-Instruct, LaMer still outperforms the RL baselines across the environments, demonstrating the general applicability of our method to different model architecture and size.

### v2

@@ Appendix A Task Description and Details
MineSweeper. We include the classic video game MineSweeper as a partially-observable environment. The game is a 2D square board, with several mines randomly scattered in the board cells. The goal of the game is to open all the safe cells without revealing the hidden mines. In each step, the agent opens a cell, and the first step is always safe. If a mine is revealed, the task ends in failure immediately. The state of the opened (safe) cells can either be empty or a number from 1 to 8, and the number specifies how many mines are adjacent to the specific cell. The agent needs to use the numbers marked on the opened cells to reason about the position of mines. Success is achieved when all safe cells are revealed. Our implementation is based on a simplified version of Li et al. (2024). The difficulty of this task is controlled by the board size and the number of mines. We train on a board size $6\times 6$ with $3$ hidden mines.
Webshop (Yao et al., 2022). We include Webshop as a partially-observable text-based environment that simulates online shopping. The agent is given a natural language instruction specifying a product to purchase with certain attributes. The environment presents a simplified e-commerce interface where the agent can search for products, navigate through search results, and examine product pages with details like price, color, size, and customer reviews. The agent must interpret the instruction, search effectively, filter through multiple product options, and select the item that best matches the specified criteria. Success is measured by whether the final purchased item satisfies all the requirements in the original instruction.
ALFWorld (Shridhar et al., 2021). We include ALFWorld as a partially-observable text-based environment that simulates household tasks in interactive fiction format. The agent receives natural language instructions for common household activities. The environment provides text descriptions of rooms, objects, and possible actions, while the agent must navigate through a house, interact with objects, and complete multi-step tasks. Objects may need to be found, picked up, cleaned, heated, or combined with other objects to achieve the goal. The agent’s view is limited to the current room and nearby objects, requiring exploration and memory of previously visited locations. Success requires understanding the instruction, planning a sequence of actions, and executing them correctly while managing partial observability. We train on the training examples of the activities ‘Pick’, ‘Look’, ‘Clean’, ‘Heat’. We evaluate in-distribution on the test examples from the same activities, and we evaluate out-of-distribution on the test examples from ‘Cool’, ’Pick2’ activities.
@@ Appendix B Example Prompts
Similar to (Feng et al., 2025), we use <action> </action> block to indicate the final decision of the action, and we use <remark> </remark> to indicate the content of the self-reflection.
@@ Appendix C Training details
LaMer is compatible with standard policy gradient algorithms. Without specification, we use GiGPO as the default optimization algorithm. During training, the self-reflection step is also explicitly trained using the reward in the subsequent episodes. During training, we match the total number of experiences sampled for each example between RL and Meta-RL to ensure a fair comparison. Specifically, for each sample $N=3$ episodes and set group size to 8 for Meta-RL, and use a group size of 24 for standard RL training. Besides that, other hyper-parameters and configuration are kept the same between RL and Meta-RL training. We use Qwen3-4B as the base model and train it with Adam optimizer and a learning rate of 1e61e-6. For Sokoban and MineSweeper, we train the agents with a batch size of $16$ for $300$ epochs. In comparison, we use batch size of $8$ and $150$ epochs for Webshop and ALFWorld. The environment reward is set to be $10$ for successful trajectories and 0 for unsuccessful ones. We use temperature of $1.0$ during rollout and $0.7$ during evaluation. The maximum number of output tokens is set to $1024$. Our code is based on the training framework of verl (Sheng et al., 2025) and verl-agent (Feng et al., 2025).
@@ D.1 Experiments on different architecture
LaMer is designed as a general framework built on meta-RL principles and is model-agnostic. To validate this, we additionally ran experiments on Llama3.1-8B-Instruct (Grattafiori et al., 2024), showing that our method works well on models in a different model architecture and model size.
@@
In Table 4, we compare LaMer with the strongest baseline GiGPO (Feng et al., 2025), on Sokoban and MineSweeper. We show The results indicate that on Llama-3.1-8B-Instruct, LaMer still outperforms the RL baselines across the environments, demonstrating the general applicability of our method to different model architecture and size.